OpenStudy (anonymous):

factor y^2-49 and 4x^2-25 and 81x^2-1

OpenStudy (anonymous):

y=(plus or minus)7, x^2=25/4 so x = (plus or minus)5/2, x^2=1/81 so x= (plus or minus) 1/9

OpenStudy (anonymous):

thanks a bunch